# Agnostic Process Tomography

### v1

@@ 1.5 Outlook
First, can one design an efficient algorithm to map an arbitrary superoperator to its closest quantum channel? This is desirable because many of our improper agnostic learning algorithms can be made proper using such a subroutine, and many applications of agnostic process tomography (Section 1.2) require proper learning algorithms. As a significant first step towards this question, we have reduced the problem to that of proving a particular mathematical condition holds (see Section 5.7) but were unable to prove or disprove it. Solving this question would be a significant result, as it would also lead to proper learning algorithms for realizable process tomography problems, such as those in [22, 30].
@@
Second, can we improve the error bounds for agnostic process tomography of bounded-gate circuits and agnostic process tomography from agnostic state tomography? Notably, as a result of the distance inequalities we used, our learning algorithms achieve error bounds of $\mathsf{opt}^{1/4}$ and $2^{1/4}\mathsf{opt}^{1/2}$, respectively, rather than $\mathsf{opt}$. We believe these bounds can be improved to $1$-agnostic process tomography and leave this as an open problem. Also, can we avoid the quadratic increase in error when converting agnostic state tomography algorithms to those for agnostic process tomography, as in quantum process tomography [6]?
@@
Third, is it possible to develop efficient agnostic learning algorithms for juntas that match the realizable lower bounds [24, 25]? Our proposed algorithms for juntas have worse complexities than the realizable setting due to needing to estimate the weight of subsets of qubits, rather than the influence.
@@
Next, for learning low-degree quantum objects, can one obtain better complexities by using the algorithm of [30] in the agnostic setting? We show in Section 4.5 that straightforwardly utilizing the algorithm from [30] results in an inefficient algorithm with worse performance than ours. However, we believe that by proving tighter upper bounds on the norms in the BH inequalities, this could be improved.
@@
Finally, we list several broad open questions. Could different definitions of agnostic tomography lead to new applications, e.g., comparing the performances of agnostic learners for different concept classes? For instance, the definition of [49] may lend itself to this, where they define agnostic learning with respect to two classes rather than just one. Can one develop agnostic learning algorithms with respect to worst-case distance metrics, e.g., diamond distance? All results in our work only hold for average-case distances. Also, many of our proposed algorithms require preparing the Choi state of the unknown channel. Instead, can one develop agnostic learning algorithms with fewer resources? As agnostic process tomography is at least as hard as in the realizable setting, lower bounds in realizable process tomography can be transferred to the agnostic setting as well. Can one prove stronger worst-case agnostic lower bounds? Finally, are there other physically-motivated classes that admit efficient agnostic process tomography algorithms? Some interesting classes include the evolution of well-studied Hamiltonians [117, 118, 119, 120] or classes of parameterized quantum circuits [60] with specific ansatz [121, 122, 123].
@@ 2 Preliminaries
For $k\in\mathbb{N}$, we write $[k]\triangleq\{1,\dots,k\}$. Throughout, let $n$ denote the system size. Let $\mathcal{H}=(\mathbb{C}^{2})^{\otimes n}$ denote the Hilbert space on these $n$ qubits, and let $\mathcal{B}(\mathcal{H})$ denote the set of bounded linear operators on $\mathcal{H}$. Let $\mathcal{U}_{d}\subseteq\mathcal{B}(\mathcal{H})$ denote the set of $d$-dimensional unitaries. A map from $\mathcal{B}(\mathcal{H})$ to itself is known as a superoperator on $\mathcal{B}(\mathcal{H})$. A quantum channel on $\mathcal{B}(\mathcal{H})$ is a superoperator that is completely positive and trace-preserving (CPTP). For any unitary, we sometimes use a calligraphic font to denote the associated quantum channel.
@@ 2.1 Choi state
A useful tool we will need is the Choi-Jamiolkowski isomorphism [124, 125], which maps between unitaries acting on the Hilbert space $\mathcal{H}$ and pure states in $\mathcal{H}\otimes\mathcal{H}$.
@@
$|v(U)\rangle\triangleq(U\otimes I)|\Phi^{+}\rangle=\frac{1}{\sqrt{2^{n}}}\sum_{i,j=0}^{2^{n}-1}U_{i,j}\ket{i,j}.$ (2.2)
@@
Lemma 1 (Properties of Choi representations of quantum channels [126]).
@@ 2.2 Pauli spectrum of operators
A key technique we use throughout our proofs is the analysis of the Pauli spectrum of quantum operators, first introduced in [28] to generalize Boolean Fourier analysis [95] to the quantum setting. Pauli analysis was also studied in [29, 24, 22]. Throughout this work, we mostly follow the notation of [28, 95].
@@
where we use $A_{i,j}$ to denote the $ij$-th entry of $A$ as a matrix. We use $\norm{\cdot}_{F}$ to denote the norm induced by this inner product, namely the Hilbert-Schmidt or Frobenius norm. In particular, we can write
$\norm{A}_{F}^{2}=\langle A,A\rangle=\sum_{i,j\in[2^{n}]}|A_{i,j}|^{2}.$ (2.7)
@@
$\sigma_{0}=I=\begin{pmatrix}1&0\\ 0&1\end{pmatrix},\quad\sigma_{1}=X=\begin{pmatrix}0&1\\ 1&0\end{pmatrix},\quad\sigma_{2}=Y=\begin{pmatrix}0&-i\\ i&0\end{pmatrix},\quad\sigma_{3}=Z=\begin{pmatrix}1&0\\ 0&-1\end{pmatrix}.$ (2.8)
For a vector $x\in\{0,1,2,3\}^{n}$, we denote $\sigma_{x}\triangleq\sigma_{x_{1}}\otimes\cdots\otimes\sigma_{x_{n}}$ and write $\mathsf{supp}(x)\triangleq\{i\in[n]:x_{i}\neq 0\}$. We will denote the degree of $x$ by $|x|=|\mathsf{supp}(x)|$. Because the Paulis $\{\sigma_{x}\}_{x\in\{0,1,2,3\}^{n}}$ form an orthogonal basis for $\mathcal{B}(\mathcal{H})$ with respect to the Hilbert-Schmidt inner product, then we can uniquely decompose any operator with respect to this basis.
Definition 4 (Pauli spectrum [28]).
@@
$A=\sum_{x\in\{0,1,2,3\}^{n}}\hat{A}_{x}\sigma_{x},\quad\hat{A}_{x}\triangleq\frac{1}{2^{n}}\langle\sigma_{x},A\rangle.$ (2.9)
@@
Lemma 2 (Proposition 10 in [28]).
@@
Following the notation in [28, 22], we define the Pauli weight of an operator at level $k$. This is similar to the classical definition of Fourier weight [95].
Definition 5 (Pauli weight at level $k$ [28, 22]).
@@
$\mathcal{P}_{S}\triangleq\{x\in\{0,1,2,3\}^{n}:\mathsf{supp}(x)\subseteq S\}.$ (2.13)
@@
This definition is reminiscent of the influence on a set of qubits. Recall from [24] that the influence of a unitary $U$ on a subset $S\subseteq[n]$ is defined as
@@
where |$\ket{v(\sigma_{x})}$ denotes the Choi state of the Pauli $\sigma_{x}$.
@@ 2.3 Fourier spectrum of superoperators
One can also extend the Pauli analysis from the previous section to superoperators, which was initiated in [25]. This is also explored in [22], but we mainly follow the presentation of [25]. The space of superoperators on $\mathcal{B}(\mathcal{H})$ can be viewed as a Hilbert space when endowed with the inner product
@@
$\Phi_{x,y}(\rho)=\sigma_{x}\rho\sigma_{y},\quad\forall x,y\in\mathcal{P}_{n}.$ (2.20)
@@
$\|\Phi_{x,y}\|_{F}=2^{n},\quad\forall x,y\in\mathcal{P}_{n}.$ (2.21)
@@
Definition 7 (Fourier expansion of superoperators; Definition 7 in [25]).
@@
$\Phi=\sum_{x,y\in\mathcal{P}_{n}}\hat{\Phi}(x,y)\Phi_{x,y},\quad\hat{\Phi}(x,y)=\frac{1}{4^{n}}\langle\Phi_{x,y},\Phi\rangle.$ (2.22)
@@
Lemma 3 (Properties of Fourier coefficients; Lemma 8 in [25]).
@@
    Then, $\hat{\Phi}\succcurlyeq 0$ if and only if $J(\Phi)\succcurlyeq 0$
@@ 2.4 Distance metrics and covering nets
Definition 10 (Normalized Frobenius distance between superoperators [25]).
@@
Lemma 6 (Normalized Frobenius distance of superoperators and Fourier coefficients; Corollary 28 in [25]).
@@
$\mathrm{d_{tr}}(\rho,\sigma)\triangleq\frac{1}{2}\norm{\rho-\sigma}_{1}.$ (2.46)
@@
We call $1-F$ the infidelity of two states.
The trace distance and fidelity are related in the following ways [127]. For any two quantum states $\rho,\sigma$,
@@
When one of the states is pure, e.g., $\sigma=\outerproduct{\psi}{\psi}$ |, then
@@
In addition to the normalized Frobenius distance, we can also consider average-case distances. We follow [27] to define an average distance between two quantum channels with respect to a measure.
Definition 12 (Average distance [27]).
@@
However, because Haar-random states are difficult to generate, one may turn to other notions of random states with respect to which to measure an average-case distance. One such physically motivated class of states is locally scrambled ensembles [128, 129, 27].
We will make use of the fact that for any measure $\mu$ over states, $d_{\mu}$ satisfies the triangle inequality. We will follow the proof of [27], where this was shown for unitaries.
Lemma 7 (Triangle inequality of $\mathrm{d_{avg}}$; Lemma 6 in [27]).
@@
By properties of the Haar integral (e.g., Example 48 in [130]), we have
@@
A useful result about covering nets that we will need in our proofs is bounds on the metric entropy of covering nets for unitaries of bounded gate complexity [27].
Lemma 9 (Covering nets for unitaries of bounded gates with respect to $d_{F}$; Theorem 8 and Corollary 1 in [27]).
@@ 3 Useful Lemmas
In this section, we compile several lemmas that we use repeatedly throughout our proofs. Our algorithms for both agnostic unitary estimation and agnostic process tomography rely on sampling from a probability distribution to estimate it empirically. In our agnostic unitary estimation algorithms, we will often sample from the distribution $\{|\hat{U}_{x}|^{2}\}_{x}$ formed by the Pauli coefficients to estimate it empirically. On the other hand, for APT, we will sample from the distribution formed by the diagonal Fourier coefficients $\{\hat{\Phi}(x,x)\}_{x}$. For guarantees on the estimation, we make use of the following well-known result from distribution learning theory. We refer to Theorem 9 in [131] for a proof and follow the presentation of Lemma 16 in [30].
Lemma 10 (Empirical estimation of a distribution; Theorem 9 in [131]).
@@
Many of our agnostic unitary estimation algorithms require learning a subset of the Pauli coefficients of a unitary. In the following lemma, we show how any Pauli coefficient can be estimated. This algorithm extends the technique introduced in Lemma 24 of [28] for learning Hermitian unitary operators, where the Pauli coefficients are all guaranteed to be real. However, in general, the Pauli coefficients may be complex. We resolve this by using the approach from [28] to separately estimate the real and imaginary parts of each Pauli coefficient. This extension is also proposed in Lemma 17 of [30] without proof. For completeness, we provide a proof here.
Lemma 11 (Estimating any Pauli coefficient of a unitary; Lemma 17 in [30]).
@@
As discussed above, the algorithm estimates the real and imaginary parts of each Pauli coefficient separately by the same approach as in Lemma 24 of [28]. To do so, we need access to controlled-$U$. This can be implemented using query access to $U$ alone by preparing $n$ ancilla qubits, applying a controlled-SWAP between the ancilla and main register, applying $U$ to the ancillas, another controlled-SWAP, and finally discarding the ancillas.
First, we estimate the real part. We proceed in the same way as Lemma 24 of [28]. We first prepare |$\ket{0}\ket{\Phi^{+}}$, where |$\ket{\Phi^{+}}$ denotes a Bell pair. Then, we apply a Hadamard gate to get
@@
Using Hoeffding’s inequality, we see that by repeating this procedure $\mathcal{O}(\log(1/\delta)/\epsilon^{2})$ many times, we can estimate $\mathrm{Re}(\hat{U}_{x})$ up to accuracy $\epsilon/\sqrt{2}$ with probability at least $21-\delta/$ 2.
@@
Then, by repeating this procedure $\mathcal{O}(\log(1/\delta)/\epsilon^{2})$ many times, we can now estimate $\mathrm{Im}(\hat{U}_{x})$ up to accuracy $\epsilon/\sqrt{2}$ with probability at least $21-\delta/$ 2. Using a union bound, we see that with probability at least $1-\delta$, we have an $\epsilon$-accurate estimate of $\hat{U}_{x}$.
@@
Similarly, we can also learn any Fourier coefficient of a quantum channel. The same result is in Lemma 15 in [30], but we give a different proof that allows for an easier analysis of the computational time. This is because our proof makes the state preparation and measurement steps more explicit than that of [30].
@@
Then, similarly to Lemma 11, by conducting $\mathcal{O}(\log(1/\delta)/\epsilon^{2})$ SWAP tests [132, 133, 134] between $v(\Phi)$ and $|v(\sigma_{x})\rangle\langle v(\sigma_{x})|$, we can estimate $\hat{\Phi}(x,x)$ up-to accuracy $\epsilon/(2\sqrt{2})$ with probability at least $41-\delta/$ 4, and similarly for $\hat{\Phi}(y,y)$. Then, consider the state
@@
where the last line uses the fact that $\hat{\Phi}(y,x)=\hat{\Phi}(x,y)^{*}$, as $\hat{\Phi}$ is Hermitian. Now, with probability at least $41-\delta/$ 4, we estimate the overlap between $v(\Phi)$ and $|\psi(x,y)^{+}\rangle\langle\psi(x,y)^{+}|$ up to accuracy $\epsilon/(2\sqrt{2})$, using $\mathcal{O}(\log(1/\delta)/\epsilon^{2})$ SWAP tests. This allows us to estimate $\mathrm{Re}(\hat{\Phi}(x,y))$ up to accuracy $\epsilon/\sqrt{2}$ since we already estimated $\hat{\Phi}(x,x)$ and $\hat{\Phi}(y,y)$ from the previous step.
@@
Another tool we utilize often is the classical shadow formalism [111], which allows for the efficient prediction of expectation values of many observables from copies of a particular state. We use classical shadows to estimate Pauli coefficients as a subroutine in some of our learning algorithms.
Lemma 13 (Classical Shadows; Theorem 1 in [111]).
@@
Finally, the last tool we need is used to ensure that the output of our agnostic learning algorithms is unitary or CPTP. This is because many of our learning algorithms rely on estimating Pauli coefficients; hence these estimates may not result in an actual unitary or quantum channel. Thus, in order to achieve proper learning, in which the output of our learning algorithm lies in the concept class $\mathcal{C}$ that we want to learn with respect to, we need to approximate our learned operator by the nearest unitary or quantum channel. We refer to this as “projection.” The following lemma shows how to achieve this for unitary projection via the approach of Lemma 8 in [23]. Note that Lemma 8 in [23] uses the same algorithm but shows that it is a unitary projection with respect to the operator norm. We show that this also gives a unitary projection with respect to the (normalized) Frobenius distance.
@@
Intuitively, the projection computes the singular value decomposition of $A$ and replaces all the singular values with $1$. This technique was originally proposed in [23], where the authors showed that this is the closest unitary projection for any operator with respect to the operator norm. We instead show that this also gives a unitary projection for the (normalized) Frobenius distance. It suffices to show that for all $B\in\mathcal{U}_{d}$
@@
concluding the proof. The runtime can be seen as the runtime of any algorithm computing the SVD and multiplying $U,V^{\dagger}$. We refer to Lecture 31 of [135] for the analysis of algorithms based on the Golub-Kahan bidiagonalization procedure [136]. Namely, computing the SVD takes time $\mathcal{O}(m^{3})$ for an $m\times m$ matrix. While such algorithms are only approximate, they demonstrate superlinear convergence, allowing one to efficiently obtain an output within machine precision and thus ignore this error in practice. ∎

### v2

@@ 1.5 Outlook
First, can one design an efficient algorithm to map an arbitrary superoperator to its closest quantum channel? This is desirable because many of our improper agnostic learning algorithms can be made proper using such a subroutine, and many applications of agnostic process tomography (Section 1.2) require proper learning algorithms. As a significant first step towards this question, we have reduced the problem to that of proving a particular mathematical condition holds (see Section 5.7) but were unable to prove or disprove it. Solving this question would be a significant result, as it would also lead to proper learning algorithms for realizable process tomography problems, such as those in [nadimpalli2024pauli, arunachalam2024learning].
@@
Second, can we improve the error bounds for agnostic process tomography of bounded-gate circuits and agnostic process tomography from agnostic state tomography? Notably, as a result of the distance inequalities we used, our learning algorithms achieve error bounds of $\mathsf{opt}^{1/4}$ and $2^{1/4}\mathsf{opt}^{1/2}$, respectively, rather than $\mathsf{opt}$. We believe these bounds can be improved to $1$-agnostic process tomography and leave this as an open problem. Also, can we avoid the quadratic increase in error when converting agnostic state tomography algorithms to those for agnostic process tomography, as in quantum process tomography [haah2023query]?
@@
Third, is it possible to develop efficient agnostic learning algorithms for juntas that match the realizable lower bounds [chen2023testing, bao2023testing]? Our proposed algorithms for juntas have worse complexities than the realizable setting due to needing to estimate the weight of subsets of qubits, rather than the influence.
@@
Next, for learning low-degree quantum objects, can one obtain better complexities by using the algorithm of [arunachalam2024learning] in the agnostic setting? We show in Section 4.5 that straightforwardly utilizing the algorithm from [arunachalam2024learning] results in an inefficient algorithm with worse performance than ours. However, we believe that by proving tighter upper bounds on the norms in the BH inequalities, this could be improved.
@@
Finally, we list several broad open questions. Could different definitions of agnostic tomography lead to new applications, e.g., comparing the performances of agnostic learners for different concept classes? For instance, the definition of [caro2023classical] may lend itself to this, where they define agnostic learning with respect to two classes rather than just one. Can one develop agnostic learning algorithms with respect to worst-case distance metrics, e.g., diamond distance? All results in our work only hold for average-case distances. Also, many of our proposed algorithms require preparing the Choi state of the unknown channel. Instead, can one develop agnostic learning algorithms with fewer resources? As agnostic process tomography is at least as hard as in the realizable setting, lower bounds in realizable process tomography can be transferred to the agnostic setting as well. Can one prove stronger worst-case agnostic lower bounds? Finally, are there other physically-motivated classes that admit efficient agnostic process tomography algorithms? Some interesting classes include the evolution of well-studied Hamiltonians [clinton2021hamiltonian, schmitt2022quantum, yanay2020two, chatterjee2024comprehensive] or classes of parameterized quantum circuits [cerezo2021variational] with specific ansatz [romero2017quantum, havlivcek2019supervised, abbas2021power].
@@ 2 Preliminaries
For $k\in\mathbb{N}$, we write $k]\triangleq\{1,\dots,k\}$. Throughout, let $n$ denote the system size. Let $\mathcal{H}=(\mathbb{C}^{2})^{\otimes n}$ denote the Hilbert space on these $n$ qubits, and let $\mathcal{B}(\mathcal{H})$ denote the set of bounded linear operators on $\mathcal{H}$. Let $\mathcal{U}_{d}\subseteq\mathcal{B}(\mathcal{H})$ denote the set of $d$-dimensional unitaries. A map from $\mathcal{B}(\mathcal{H})$ to itself is known as a superoperator on $\mathcal{B}(\mathcal{H})$. A quantum channel on $\mathcal{B}(\mathcal{H})$ is a superoperator that is completely positive and trace-preserving (CPTP). For any unitary, we sometimes use a calligraphic font to denote the associated quantum channel.
@@ 2.1 Choi state
A useful tool we will need is the Choi-Jamiolkowski isomorphism [choi1975completely, jamiolkowski1972linear], which maps between unitaries acting on the Hilbert space $\mathcal{H}$ and pure states in $\mathcal{H}\otimes\mathcal{H}$.
@@
[…]
@@
Lemma 1 (Properties of Choi representations of quantum channels [watrous2018theory]).
@@ 2.2 Pauli spectrum of operators
A key technique we use throughout our proofs is the analysis of the Pauli spectrum of quantum operators, first introduced in [montanaro2008quantum] to generalize Boolean Fourier analysis [o2014analysis] to the quantum setting. Pauli analysis was also studied in [rouze2024quantum, chen2023testing, nadimpalli2024pauli]. Throughout this work, we mostly follow the notation of [montanaro2008quantum, o2014analysis].
@@
where we use $A_{i,j}$ to denote the $ij$-th entry of $A$ as a matrix. We use $\mathinner{\!\left\lVert\cdot\right\rVert}_{F}$ to denote the norm induced by this inner product, namely the Hilbert-Schmidt or Frobenius norm. In particular, we can write
$\mathinner{\!\left\lVert A\right\rVert}_{F}^{2}=\langle A,A\rangle=\sum_{i,j\in[2^{n}]}|A_{i,j}|^{2}.$ (2.7)
@@
$\sigma_{0}=I=\begin{pmatrix}1&0\\ 0&1\end{pmatrix},\qquad\sigma_{1}=X=\begin{pmatrix}0&1\\ 1&0\end{pmatrix},\qquad\sigma_{2}=Y=\begin{pmatrix}0&-i\\ i&0\end{pmatrix},\qquad\sigma_{3}=Z=\begin{pmatrix}1&0\\ 0&-1\end{pmatrix}$. (2.8)
For a vector $x\in\{0,1,2,3\}^{n}$, we denote $\sigma_{x}\triangleq\sigma_{x_{1}}\otimes\cdots\otimes\sigma_{x_{n}}$ and write $\mathsf{supp}(x)\triangleq\{i\in[n]\mathrel{\mathop{\ordinarycolon}}x_{i}\neq 0\}$. We will denote the degree of $x$ by $|x|=|\mathsf{supp}(x)|$. Because the Paulis $\{\sigma_{x}\}_{x\in\{0,1,2,3\}^{n}}$ form an orthogonal basis for $\mathcal{B}(\mathcal{H})$ with respect to the Hilbert-Schmidt inner product, then we can uniquely decompose any operator with respect to this basis.
Definition 4 (Pauli spectrum [montanaro2008quantum]).
@@
$A=\sum_{x\in\{0,1,2,3\}^{n}}\hat{A}_{x}\sigma_{x},\qquad\hat{A}_{x}\triangleq\frac{1}{2^{n}}\langle\sigma_{x},A\rangle.$ (2.9)
@@
Lemma 2 (Proposition 10 in [montanaro2008quantum]).
@@
Following the notation in [montanaro2008quantum, nadimpalli2024pauli], we define the Pauli weight of an operator at level $k$. This is similar to the classical definition of Fourier weight [o2014analysis].
Definition 5 (Pauli weight at level $k$ [montanaro2008quantum, nadimpalli2024pauli]).
@@
$\mathcal{P}_{S}\triangleq\{x\in\{0,1,2,3\}^{n}\mathrel{\mathop{\ordinarycolon}}\mathsf{supp}(x)\subseteq S\}.$ (2.13)
@@
This definition is reminiscent of the influence on a set of qubits. Recall from [chen2023testing] that the influence of a unitary $U$ on a subset $S\subseteq[n]$ is defined as
@@
[…]
@@ 2.3 Fourier spectrum of superoperators
One can also extend the Pauli analysis from the previous section to superoperators, which was initiated in [bao2023testing]. This is also explored in [nadimpalli2024pauli], but we mainly follow the presentation of [bao2023testing]. The space of superoperators on $\mathcal{B}(\mathcal{H})$ can be viewed as a Hilbert space when endowed with the inner product
@@
$\Phi_{x,y}(\rho)=\sigma_{x}\rho\sigma_{y},\qquad\forall x,y\in\mathcal{P}_{n}.$ (2.20)
@@
$\|\Phi_{x,y}\|_{F}=2^{n},\qquad\forall x,y\in\mathcal{P}_{n}.$ (2.21)
@@
Definition 7 (Fourier expansion of superoperators; Definition 7 in [bao2023testing]).
@@
$\Phi=\sum_{x,y\in\mathcal{P}_{n}}\hat{\Phi}(x,y)\Phi_{x,y},\qquad\hat{\Phi}(x,y)=\frac{1}{4^{n}}\langle\Phi_{x,y},\Phi\rangle.$ (2.22)
@@
Lemma 3 (Properties of Fourier coefficients; Lemma 8 in [bao2023testing]).
@@
    Then, $\hat{\Phi}\succcurlyeq 0$ if and only if $J(\Phi)\succcurlyeq 0$ .
@@ 2.4 Distance metrics and covering nets
Definition 10 (Normalized Frobenius distance between superoperators [bao2023testing]).
@@
Lemma 6 (Normalized Frobenius distance of superoperators and Fourier coefficients; Corollary 28 in [bao2023testing]).
@@
$\mathrm{d_{tr}}(\rho,\sigma)\triangleq\frac{1}{2}\mathinner{\!\left\lVert\rho-\sigma\right\rVert}_{1}.$ (2.46)
@@
We call 1F1-F the infidelity of two states.
The trace distance and fidelity are related in the following ways [nielsen2010quantum]. For any two quantum states $\rho,\sigma$,
@@
When one of the states is pure, e.g., $|\sigma=\outerproduct{\psi}{\psi}$, then
@@
In addition to the normalized Frobenius distance, we can also consider average-case distances. We follow [zhao2023learning] to define an average distance between two quantum channels with respect to a measure.
Definition 12 (Average distance [zhao2023learning]).
@@
However, because Haar-random states are difficult to generate, one may turn to other notions of random states with respect to which to measure an average-case distance. One such physically motivated class of states is locally scrambled ensembles [kuo2020markovian, hu2023classical, zhao2023learning].
We will make use of the fact that for any measure $\mu$ over states, $d_{\mu}$ satisfies the triangle inequality. We will follow the proof of [zhao2023learning], where this was shown for unitaries.
Lemma 7 (Triangle inequality of $\mathrm{d_{avg}}$; Lemma 6 in [zhao2023learning]).
@@
By properties of the Haar integral (e.g., Example 48 in [mele2024introduction]), we have
@@
A useful result about covering nets that we will need in our proofs is bounds on the metric entropy of covering nets for unitaries of bounded gate complexity [zhao2023learning].
Lemma 9 (Covering nets for unitaries of bounded gates with respect to $d_{F}$; Theorem 8 and Corollary 1 in [zhao2023learning]).
@@ 3 Useful Lemmas
In this section, we compile several lemmas that we use repeatedly throughout our proofs. Our algorithms for both agnostic unitary estimation and agnostic process tomography rely on sampling from a probability distribution to estimate it empirically. In our agnostic unitary estimation algorithms, we will often sample from the distribution $\{|\hat{U}_{x}|^{2}\}_{x}$ formed by the Pauli coefficients to estimate it empirically. On the other hand, for APT, we will sample from the distribution formed by the diagonal Fourier coefficients $\{\hat{\Phi}(x,x)\}_{x}$. For guarantees on the estimation, we make use of the following well-known result from distribution learning theory. We refer to Theorem 9 in [canonne2020short] for a proof and follow the presentation of Lemma 16 in [arunachalam2024learning].
Lemma 10 (Empirical estimation of a distribution; Theorem 9 in [canonne2020short]).
@@
Many of our agnostic unitary estimation algorithms require learning a subset of the Pauli coefficients of a unitary. In the following lemma, we show how any Pauli coefficient can be estimated. This algorithm extends the technique introduced in Lemma 24 of [montanaro2008quantum] for learning Hermitian unitary operators, where the Pauli coefficients are all guaranteed to be real. However, in general, the Pauli coefficients may be complex. We resolve this by using the approach from [montanaro2008quantum] to separately estimate the real and imaginary parts of each Pauli coefficient. This extension is also proposed in Lemma 17 of [arunachalam2024learning] without proof. For completeness, we provide a proof here.
Lemma 11 (Estimating any Pauli coefficient of a unitary; Lemma 17 in [arunachalam2024learning]).
@@
As discussed above, the algorithm estimates the real and imaginary parts of each Pauli coefficient separately by the same approach as in Lemma 24 of [montanaro2008quantum]. To do so, we need access to controlled-$U$. This can be implemented using query access to $U$ alone by preparing $n$ ancilla qubits, applying a controlled-SWAP between the ancilla and main register, applying $U$ to the ancillas, another controlled-SWAP, and finally discarding the ancillas.
@@
Using Hoeffding’s inequality, we see that by repeating this procedure $\mathcal{O}(\log(1/\delta)/\epsilon^{2})$ many times, we can estimate $\mathrm{Re}(\hat{U}_{x})$ up to accuracy $\epsilon/\sqrt{2}$ with probability at least $1-\delta/2$.
@@
Then, by repeating this procedure $\mathcal{O}(\log(1/\delta)/\epsilon^{2})$ many times, we can now estimate $\mathrm{Im}(\hat{U}_{x})$ up to accuracy $\epsilon/\sqrt{2}$ with probability at least $1-\delta/2$. Using a union bound, we see that with probability at least $1-\delta$, we have an $\epsilon$-accurate estimate of $\hat{U}_{x}$.
@@
Similarly, we can also learn any Fourier coefficient of a quantum channel. The same result is in Lemma 15 in [arunachalam2024learning], but we give a different proof that allows for an easier analysis of the computational time. This is because our proof makes the state preparation and measurement steps more explicit than that of [arunachalam2024learning].
@@
Then, similarly to Lemma 11, by conducting $\mathcal{O}(\log(1/\delta)/\epsilon^{2})$ SWAP tests [buhrman2001quantum, barenco1997stabilization, kobayashi2003quantum] between $v(\Phi)$ and $|v(\sigma_{x})\rangle\langle v(\sigma_{x})|$, we can estimate $\hat{\Phi}(x,x)$ up-to accuracy $\epsilon/(2\sqrt{2})$ with probability at least $1-\delta/4$, and similarly for $\hat{\Phi}(y,y)$. Then, consider the state
@@
where the last line uses the fact that $\hat{\Phi}(y,x)=\hat{\Phi}(x,y)^{*}$, as $\hat{\Phi}$ is Hermitian. Now, with probability at least $1-\delta/4$, we estimate the overlap between $v(\Phi)$ and $|\psi(x,y)^{+}\rangle\langle\psi(x,y)^{+}|$ up to accuracy $\epsilon/(2\sqrt{2})$, using $\mathcal{O}(\log(1/\delta)/\epsilon^{2})$ SWAP tests. This allows us to estimate $\mathrm{Re}(\hat{\Phi}(x,y))$ up to accuracy $\epsilon/\sqrt{2}$ since we already estimated $\hat{\Phi}(x,x)$ and $\hat{\Phi}(y,y)$ from the previous step.
@@
Another tool we utilize often is the classical shadow formalism [huang2020predicting], which allows for the efficient prediction of expectation values of many observables from copies of a particular state. We use classical shadows to estimate Pauli coefficients as a subroutine in some of our learning algorithms.
Lemma 13 (Classical Shadows; Theorem 1 in [huang2020predicting]).
@@
Finally, the last tool we need is used to ensure that the output of our agnostic learning algorithms is unitary or CPTP. This is because many of our learning algorithms rely on estimating Pauli coefficients; hence these estimates may not result in an actual unitary or quantum channel. Thus, in order to achieve proper learning, in which the output of our learning algorithm lies in the concept class $\mathcal{C}$ that we want to learn with respect to, we need to approximate our learned operator by the nearest unitary or quantum channel. We refer to this as “projection.” The following lemma shows how to achieve this for unitary projection via the approach of Lemma 8 in [huang2024learning]. Note that Lemma 8 in [huang2024learning] uses the same algorithm but shows that it is a unitary projection with respect to the operator norm. We show that this also gives a unitary projection with respect to the (normalized) Frobenius distance.
@@
Intuitively, the projection computes the singular value decomposition of $A$ and replaces all the singular values with $1$. This technique was originally proposed in [huang2024learning], where the authors showed that this is the closest unitary projection for any operator with respect to the operator norm. We instead show that this also gives a unitary projection for the (normalized) Frobenius distance. It suffices to show that for all $B\in\mathcal{U}_{d}$
@@
concluding the proof. The runtime can be seen as the runtime of any algorithm computing the SVD and multiplying $U,V^{\dagger}$. We refer to Lecture 31 of [trefethen2022numerical] for the analysis of algorithms based on the Golub-Kahan bidiagonalization procedure [golub1965calculating]. Namely, computing the SVD takes time $\mathcal{O}(m^{3})$ for an $m\times m$ matrix. While such algorithms are only approximate, they demonstrate superlinear convergence, allowing one to efficiently obtain an output within machine precision and thus ignore this error in practice. ∎